# Nonlinear Separation Theorems for Co-Radiant Sets and Optimality Conditions for Approximate and Proper Approximate Solutions in Vector Optimization

### v2

@@ 2. Notation and Previous Terminology
A subset $C\subset X$ is said to be a cone if $\lambda x\in C$ for every $\lambda\geq 0$ and $x\in C$. A cone $C$ is said to be a convex cone if $C$ is a convex subset of $X$. Fixed a subset $S\subset X$, we define the cone generated by $S$ as cone$(S):=\{\lambda s\colon\lambda\geq 0,\,s\in S\}$ and $\overline{\mbox{cone}}(S)$ stands for the closure of cone$(S)$. A non-empty convex subset $B$ of a convex cone $C$ is said to be a base of $C$ if $0_{X}\not\in\overline{B}$ and for every $x\in C$, $x\not=0_{X}$, there exists unique $\lambda_{x}>0$ and $b_{x}\in B$ such that $x=\lambda_{x}b_{x}$. A convex cone $C$ is said to have a bounded base if there exists a base $B$ of $C$ such that it is a bounded subset of $X$. Given a cone $C\subset X$, it is defined the dual cone of $C$ by $C^{*}:=\{f\in X^{*}\colon f(x)\geq 0,\,\forall x\in C\}$ and the quasi-interior of $C^{*}$ by $C^{\#}:=\{f\in X^{*}\colon f(x)>0,\,\forall x\in C\setminus\{0_{X}\}\}$. It is known that $C^{\#}\supset\mbox{int}(C^{*})$ and that any convex cone $C\subset X$ has a base if and only if $C^{\#}\not=\emptyset$.
The following sets were introduced in [4] and they are called augmented dual cones, $C^{a*}_{+}:=\{(f,\alpha)\in C^{\#}\times(0,+\infty)\colon f(x)-\alpha\|x\|\geq 0,\,\forall x\in C\}$ and $C^{a\#}_{+}:=\{(f,\alpha)\in C^{\#}\times(0,+\infty)\colon f(x)-\alpha\|x\|>0,\,\forall x\in C\setminus\{0_{X}\}\}$. A function $p:X\rightarrow{\rm I}\hskip-1.30005pt{\rm R}$ is said to be sublinear if it is both positively homogeneous and subadditive, i.e., $p(\alpha x)=\alpha p(x)$, for every $\alpha\geq 0$, and $p(x+y)\leq p(x)+p(y)$, for all $x$, $y\in X$, respectively.
@@
${{\mathcal{C}}}^{a*}_{\lambda}(\varepsilon):=\left\{(f,\alpha)\in{{\mathcal{C}}}^{\#}\times\mathbb{R}_{+}:f(y)-\alpha\parallel y\parallel\geq\lambda,\,\forall y\in{{\mathcal{C}}}(\varepsilon)\right\},\$
@@ 3. Some Preliminary Results Regarding Cones and Co-Radiant Sets
Next, we introduce (from [2]) the concept of $\delta$-conic neighborhood of a cone.
@@
Let $X$ be a normed space, let $C\subset X$ be a cone, $10<\delta<$ 1, and consider $S_{\delta}:=\{y\in X\colon d(y,C\cap S_{X})\leq\delta\}$. The $\delta$-conic neighbourhood of $C$ is the cone defined by $C_{\delta}:=\mbox{cone}(S_{\delta})$.
@@
Assume that the pair $(C,K)$ satisfies the SSP. By Theorem 3.2, there exist $\delta_{1},\delta_{2}\in\mathbb{R}$ and $f\in C^{\#}$ such that $0<\delta_{1}<\delta_{2}$, and for every $\alpha\in(\delta_{1},\delta_{2})$, we have $(f,\alpha)\in C^{a\#}_{+}$ satisfying $f(x)+\alpha\|x\|<0<f(y)+\alpha\|y\|$ for all $x\in-C\setminus\{0_{X}\}$ and $y\in\mathrm{bd}(-K)\setminus\{0_{X}\}$. Now pick $\alpha_{0},\alpha_{1},\alpha_{2}\in\mathbb{R}$ such that $0<\delta_{1}<\alpha_{1}<\alpha_{0}<\alpha_{2}<\delta_{2}$ and consider $\delta^{\prime}:=\frac{\alpha_{2}-\alpha_{0}}{2(\|f\|+\alpha_{0})}$. Let $\delta\in(0,\delta^{\prime})$ and fix $x\in-\mbox{cone}(S_{X}\cap C+2\delta B_{X})\setminus\{0_{X}\}$. Then $f(x)+\alpha\|x\|<0$ for every $\alpha\in[\alpha_{1},\alpha_{0}]$. Indeed, we write $x=-\lambda(s+2\delta b)$ for some $s\in C\cap S_{X}$, $b\in B_{X}$, and $\lambda>0$. Thus $f(x)+\alpha\|x\|\leq f(x)+\alpha_{0}\|x\|=f(-\lambda(s+2\delta b))+\alpha_{0}\|\lambda(s+2\delta b)\|=$
$=f(-\lambda s)+2\delta\lambda f(-b)+\alpha_{0}\|\lambda s+2\lambda\delta b\|\quad\text{(by the linearity of $f\in X^{*}$)}$
[…]
$\leq f(-\lambda s)+2\delta\lambda\|f\|_{*}\cdot\|b\|+\alpha_{0}\|\lambda s\|+2\alpha_{0}\lambda\delta\|b\|\quad\text{(since $f(-b)\leq|f(-b)|\leq\|f\|_{*}\cdot\|b\|$)}$
$\leq f(-\lambda s)+\alpha_{0}\|\lambda s\|+2\lambda\delta\|f\|_{*}+2\alpha_{0}\lambda\delta\quad\text{(since $\|b\|\leq 1$)}$
$=f(-\lambda s)+\alpha_{0}\|\lambda s\|+2\lambda\delta(\|f\|_{*}+\alpha_{0})\leq f(-\lambda s)+\alpha_{0}\|\lambda s\|+\lambda(\alpha_{2}-\alpha_{0})=f(-\lambda s)+\alpha_{0}\|\lambda s\|+(\alpha_{2}-\alpha_{0})\|\lambda s\|=f(-\lambda s)+\alpha_{2}\|-\lambda s\|<0$. Now, since $C_{\delta}\ =\mbox{cone}(\overline{S_{X}\cap C+\delta B_{X}})\subset\mbox{cone}(S_{X}\cap C+2\delta B_{X})$, it follows that for every $\alpha\in[\alpha_{1},\alpha_{0}]$ we have $f(x)+\alpha\|x\|<0<f(y)+\alpha\|y\|$ whenever $x\in-C_{\delta}\setminus\{0_{X}\}$ and $y\in\mbox{bd}(-K)\setminus\{0_{X}\}$. Furthermore, the inequality $f(x)+\alpha\|x\|<0$ for every $x\in-C_{\delta}\setminus\{0_{X}\}$ leads to $f\in(C_{\delta})^{a\#}_{+}$. Then, Theorem 3.2 (ii) $\Rightarrow$ (i) applies, concluding that the pair $(C_{\delta},K)$ satisfies the SSP. ∎
@@
In order to simplify the notation, we will denote $C=\mbox{cone}({{\mathcal{C}}})$ and $K_{t}=\mbox{cone}({{\mathcal{K}}}_{t})$. Let us show (i)$\Leftrightarrow$(ii) directly. By definition, the pair $(C,K_{t})$ satisfies the SSP if and only if
@@
$0_{X}\not\in t\overline{\mbox{co}(C\cap S_{X})-\mbox{co}((\mbox{bd}(K_{t})\cap S_{X})\cup\{0_{X}\}))},\mbox{ for any }t>0$
@@
Proposition 3.8 and assertion (i) in Lemma 3.7 give (i)$\Leftrightarrow$(ii). For (ii)$\Leftrightarrow$(iii) we apply Proposition 3.8 and assertions (i), (ii), and (iv) in Lemma 3.7
@@ 4. Separation Theorems for Co-Radiant Sets
[…]
@@
[…]
@@
    For every $x\in X$ we have the inequalities $-\|f\|_{*}\|x\|\leq p_{f,\alpha}(x)<2\|f\|_{*}\|x\|$.
@@
    $-\rho f(x^{\prime})+\alpha\rho<-\rho(\alpha+\beta)+\alpha\rho=-\rho\beta.$
@@
Let $C\subset X$ be a cone, take $10<\delta<$ 1, $d>0$, and $y\in C_{\geq d}$. Then $y+\delta dB_{X}\subset(C_{\delta})_{\geq(1-\delta)d}$.
@@
(i)$\Rightarrow$(ii) Assume that the pair $(\mbox{cone}({{\mathcal{C}}}),\mbox{cone}({{\mathcal{K}}}))$ satisfies the SSP and denote $K=\mbox{cone}({{\mathcal{K}}})$ and $C=\mbox{cone}({{\mathcal{C}}})$. By Lemma 3.4, there exists $10<\delta<$ 1 such that the pair $(C_{\delta},K)$ satisfies the SSP. Now, by Theorem 3.2, there exist $\alpha_{1},\alpha_{2}\in\mathbb{R}$ with $0<\alpha_{1}<\alpha_{2}$, and a functional $f\in(C_{\delta})^{\#}$ such that $(f,\alpha)\in(C_{\delta})^{a\#}_{+}$ for every $\alpha\in(\alpha_{1},\alpha_{2})$, and
@@
for every $y\in-C_{\delta}\setminus\{0_{X}\}$ and $y^{\prime}\in\mbox{bd}(-K)\setminus\{0_{X}\}$. Furthermore, $f\in{{\mathcal{C}}}^{\#}$ because $f\in(C_{\delta})^{\#}\subset C^{\#}$ and ${{\mathcal{C}}}\subset C\setminus\{0_{X}\}$. On the other hand, without loss of generality, we may assume that $f\in S_{X^{*}}$. Define $\lambda:=\inf\{f(y)-\alpha\|y\|:y\in{{\mathcal{C}}},\alpha\in(\alpha_{1},\alpha_{2})\}$, we will prove that $\lambda>0$ or, equivalently, that $-\lambda=\sup\{p_{f,\alpha}(y):y\in-{{\mathcal{C}}},\alpha\in(\alpha_{1},\alpha_{2})\}<0$. The inequality $-\lambda\leq 0$ is a direct consequence of (2) because $-{{\mathcal{C}}}\subset-C_{\delta}\setminus\{0_{X}\}$. Now assume, contrary to our claim, that $\lambda=0$. For every $\alpha\in(\alpha_{1},\alpha_{2})$, we have $(f,\alpha)\in(C_{\delta})^{a\#}_{+}\subset C^{a\#}_{+}$, which implies $\alpha<1$ as $f$ is taken in $S_{X^{*}}$. Then $\alpha_{2}\leq 1$. However, without loss of generality, we may suppose that $\alpha_{2}<1$ because otherwise we can redefine $\alpha_{2}\in(\alpha_{1},1)$. Now, consider $\mu=\frac{1-\alpha_{2}}{2}\delta d_{{{\mathcal{C}}}}>0$. By the assumption $\lambda=0$, there exist $y_{0}\in-{{\mathcal{C}}}\subset-C_{\geq d_{{{\mathcal{C}}}}}$ and $\alpha_{0}\in(\alpha_{1},\alpha_{2})$ such that
$-\mu<p_{f,\alpha_{0}}(y_{0})<0.$ (3)
@@
$-\lambda<p_{f,\alpha}(y^{\prime}),\,\forall\alpha_{1}<\alpha<\alpha_{2},\,y^{\prime}\in\mbox{bd}(-{{\mathcal{K}}}(\eta)).$ (6)
To prove (6), we fix $\xi>0$ such that $\eta=\frac{\lambda}{\xi+I_{{{\mathcal{K}}}}}$ and define $\tau:=\lambda\frac{\frac{\xi}{2}+I_{{{\mathcal{K}}}}}{\xi+I_{{{\mathcal{K}}}}}<\lambda$. By Proposition 4.2, we have $I_{{{\mathcal{K}}}(\eta)}=\eta I_{{{\mathcal{K}}}}<(\frac{\xi}{2}+I_{{{\mathcal{K}}}})\eta=\tau$. Now pick an arbitrary ${\bar{y}}\in\mbox{bd}(-{{\mathcal{K}}}(\eta))$. If $\|{\bar{y}}\|\geq\tau$, then, by Lemma 3.7 (v), ${\bar{y}}\in\mbox{bd}(\mbox{cone}(-{{\mathcal{K}}}(\eta)))=\mbox{bd}(\mbox{cone}(-{{\mathcal{K}}}))=\mbox{bd}(-{{\mathcal{K}}})$. Now (2) applies and we have $-\lambda<0<p_{f,\alpha}(\bar{y})$, for every $\alpha\in(\alpha_{1},\alpha_{2})$. To complete the proof, assume now that $\|{{\bar{y}}}\|<\tau$. Then Lemma 4.3 (iii) assures that $-\lambda<-\tau\leq p_{f,\alpha}(\bar{y})$, for every $\alpha\in(\alpha_{1},\alpha_{2})$ finishing the proof of (6).
(ii)$\Rightarrow$(i) By assumption, we consider $0<\alpha_{1}<\alpha_{2}$ and $f\in S_{X^{*}}$ such that $\lambda:=\inf\{f(y)-\alpha\|y\|:y\in{{\mathcal{C}}},\alpha\in(\alpha_{1},\alpha_{2})\}>0$ and, if $0<\eta<\frac{\lambda}{I_{{{\mathcal{K}}}}}$, then $f(y^{\prime})-\alpha\|y^{\prime}\|<\lambda<f(y)-\alpha\|y\|$, for every $\alpha_{1}<\alpha<\alpha_{2}$, $y^{\prime}\in\mbox{bd}({{\mathcal{K}}}(\eta))$, and $y\in\overline{\mbox{co}}({{\mathcal{C}}})$. Now, assume that ${{\mathcal{C}}}$ admits a norm base. We fix $t>\max\{I_{{{\mathcal{K}}}(\eta)},I_{{{\mathcal{C}}}}\}$, note that $t>0$ because $I_{{{\mathcal{K}}}(\eta)}=\eta I_{{{\mathcal{K}}}}>0$ by Proposition 4.2. For any $y\in{{\mathcal{C}}}\cap tS_{X}$ and $y^{\prime}\in(\mbox{bd}({{\mathcal{K}}}(\eta))\cap tS_{X})\cup\{0_{X}\}$ we have that $f(y)>\lambda+\alpha t>f(y^{\prime})$ for all $\alpha\in(\alpha_{1},\alpha_{2})$. Now, fix $\alpha_{3},\alpha_{4}\in\mathbb{R}$ satisfying $\alpha_{1}<\alpha_{3}<\alpha_{4}<\alpha_{2}$. Then, we have $f(y)>\lambda+\alpha_{4}t>\lambda+\alpha_{3}t>f(y^{\prime})$, for every $y\in{{\mathcal{C}}}\cap tS_{X}$ and $y^{\prime}\in(\mbox{bd}({{\mathcal{K}}}(\eta))\cap tS_{X})\cup\{0_{X}\}$, which also hold for $y\in\mbox{co}({{\mathcal{C}}}\cap tS_{X})$ and $y^{\prime}\in\mbox{co}((\mbox{bd}({{\mathcal{K}}}(\eta))\cap tS_{X})\cup\{0_{X}\})$. As a consequence, we have the inequality
@@
Next, we state our second separation theorem. In this statement, there is one more assumption than in the previous one; specifically, it is assumed that the co-radiant sets fit into each other. This hypothesis allows for the separation of the set $\overline{\mbox{co}}({{\mathcal{C}}})$ and the complementary of the set $\mbox{int}(K(\eta))$ by the same separating surface defined in Theorem 4.1. Regarding the proof, we only need to justify that (i)$\Rightarrow$(ii), but such a proof is the same as in the implication (i)$\Rightarrow$(ii) of Theorem 4.1, but using [4, Theorem 3.3] instead of Theorem 3.2.
@@
Then (i)$\Rightarrow$(ii). Moreover, if ${{\mathcal{C}}}$ also admits a norm base, then (ii)$\Rightarrow$(i).
@@
    ${{\mathcal{K}}}_{3}:=\{x\in\mathbb{R}^{2}:f(x)-\alpha\|x\|_{2}\geq\lambda\}$, for any $f\in(\mathbb{R}^{2})^{*}$, $0<\alpha<\|f\|_{2}$, and $\lambda>0$, being $\|\cdot\|_{2}$ the euclidean norm on $\mathbb{R}^{2}$.
Despite being uncommon for a co-radiant set to admit a norm base, what does occur is that any co-radiant set can be expressed as a union of a monotonically non-decreasing sequence (in the sense of inclusion) of co-radiant subsets with a norm base. Indeed, let ${{\mathcal{K}}}\subset X$ be a co-radiant set. Let us recall that, for every $r>0$, we define the set ${{\mathcal{K}}}_{r}:={{\mathcal{K}}}\cap rS_{X}$. Now, for every $n\in\mathbb{N}$ we define ${{\mathcal{K}}}^{n}:={{\mathcal{K}}}\cap\mbox{cone}({{\mathcal{K}}}_{n})$, if ${{\mathcal{K}}}_{n}\not=\emptyset$, and ${{\mathcal{K}}}^{n}=\emptyset$, otherwise. The following lemma is easy to check.
@@
    If ${{\mathcal{K}}}_{n}\not=\emptyset$, then ${{\mathcal{K}}}_{n}={{\mathcal{K}}}^{n}\cap nS_{X}$ and ${{\mathcal{K}}}_{n}$ is a norm-base for ${{\mathcal{K}}}^{n}$.
@@
Let $X$ be a normed space, and let ${{\mathcal{C}}}$ and ${{\mathcal{K}}}$ be co-radiant subsets of $X$ such that $d_{{{\mathcal{C}}}}d_{{{\mathcal{K}}}}>0$ and $\overline{co}({{\mathcal{C}}})\cap{{\mathcal{K}}}\not=\varnothing$. Assume that there exists $n\in\mathbb{N}$ such that the pair $(\mbox{cone}({{\mathcal{C}}}),\mbox{cone}({{\mathcal{K}}}_{n}))$ satisfies the SSP. Then, there exist $\alpha_{1}$, $\alpha_{2}\in\mathbb{R}$ with $0<\alpha_{1}<\alpha_{2}$ and $f\in S_{X^{*}}$ such that:
@@
    $f(y^{\prime})-\alpha\|y^{\prime}\|<\lambda<f(y)-\alpha\|y\|,$ (9)

### v3

@@ 2. Notation and Previous Terminology
Throughout the paper, let $X$ denote a normed space, with $\|\cdot\|$ denoting its norm, $X^{*}$ the dual Banach space of $X$, and $\|\cdot\|_{*}$ the corresponding dual norm. We denote by $B_{X}:=\{x\in X:\|x\|\leq 1\}$ (resp. $B_{X^{*}}:=\{f\in X^{*}:\|f\|_{*}\leq 1\}$) the closed unit ball of $X$ (resp. of $X^{*}$), and by $S_{X}:=\{x\in X:\|x\|=1\}$ (resp. $S_{X^{*}}:=\{f\in X^{*}:\|f\|_{*}=1\}$) the unit sphere of $X$ (resp. of $X^{*}$). Additionally, we define $B_{X}(x,\varepsilon):=\{x^{\prime}\in X:\|x-x^{\prime}\|\leq\varepsilon\}$.
@@
The following sets were introduced in [4] and they are called augmented dual cones, $C^{a*}_{+}:=\{(f,\alpha)\in C^{\#}\times(0,+\infty)\colon f(x)-\alpha\|x\|\geq 0,\,\forall x\in C\}$ and $C^{a\#}_{+}:=\{(f,\alpha)\in C^{\#}\times(0,+\infty)\colon f(x)-\alpha\|x\|>0,\,\forall x\in C\setminus\{0_{X}\}\}$. A function $p:X\rightarrow{\rm I}\thinspace{\rm R}$ is said to be sublinear if it is both positively homogeneous and subadditive, i.e., $p(\alpha x)=\alpha p(x)$, for every $\alpha\geq 0$, and $p(x+y)\leq p(x)+p(y)$, for all $x$, $y\in X$, respectively.
@@
${{\mathcal{C}}}^{a*}_{\lambda}(\varepsilon):=\left\{(f,\alpha)\in{{\mathcal{C}}}^{\#}\times\mathbb{R}_{+}:f(y)-\alpha\parallel y\parallel\geq\lambda,\,\forall y\in{{\mathcal{C}}}(\varepsilon)\right\},\\$
@@ 3. Some Preliminary Results Regarding Cones and Co-Radiant Sets
Next, we introduce (from [2]) the concept of a $\delta$-conic neighborhood of a cone.
@@
Let $X$ be a normed space, let $C\subset X$ be a cone, and let $0<\delta<1$. For any $y\in X$ and set $A\subset X$, let $d(y,A):=\inf_{a\in A}\|y-a\|$ denote the distance from $y$ to $A$. Consider $S_{\delta}:=\{y\in X\colon d(y,C\cap S_{X})\leq\delta\}$. The $\delta$-conic neighbourhood of $C$ is the cone defined by $C_{\delta}:=\mathrm{cone}(S_{\delta})$.
@@
Assume that the pair $(C,K)$ satisfies the SSP. By Theorem 3.2, there exist $\delta_{1},\delta_{2}\in\mathbb{R}$ and $f\in C^{\#}$ such that $0<\delta_{1}<\delta_{2}$, and for every $\alpha\in(\delta_{1},\delta_{2})$, we have $(f,\alpha)\in C^{a\#}_{+}$ satisfying $f(x)+\alpha\|x\|<0<f(y)+\alpha\|y\|$ for all $x\in-C\setminus\{0_{X}\}$ and $y\in\mathrm{bd}(-K)\setminus\{0_{X}\}$. Now pick $\alpha_{0},\alpha_{1},\alpha_{2}\in\mathbb{R}$ such that $0<\delta_{1}<\alpha_{1}<\alpha_{0}<\alpha_{2}<\delta_{2}$ and consider $\delta^{\prime}:=\frac{\alpha_{2}-\alpha_{0}}{2(\|f\|+\alpha_{0})}$. Let $\delta\in(0,\delta^{\prime})$ and fix $x\in-\mbox{cone}(S_{X}\cap C+2\delta B_{X})\setminus\{0_{X}\}$. Then $f(x)+\alpha\|x\|<0$ for every $\alpha\in[\alpha_{1},\alpha_{0}]$. Indeed, we write $x=-\lambda(s+2\delta b)$ for some $s\in C\cap S_{X}$, $b\in B_{X}$, and $\lambda>0$. Thus
$f(x)+\alpha\|x\|$ $\leq f(x)+\alpha_{0}\|x\|$
$=f(-\lambda(s+2\delta b))+\alpha_{0}\|\lambda(s+2\delta b)\|$
$=f(-\lambda s)+2\delta\lambda f(-b)+\alpha_{0}\|\lambda s+2\lambda\delta b\|$
$\leq f(-\lambda s)+2\delta\lambda f(-b)+\alpha_{0}\|\lambda s\|+2\alpha_{0}\lambda\delta\|b\|$
$\leq f(-\lambda s)+2\delta\lambda\|f\|_{*}\|b\|+\alpha_{0}\|\lambda s\|+2\alpha_{0}\lambda\delta\|b\|$
$\leq f(-\lambda s)+2\delta\lambda\|f\|_{*}+\alpha_{0}\|\lambda s\|+2\alpha_{0}\lambda\delta$
$=f(-\lambda s)+\alpha_{0}\|\lambda s\|+2\delta\lambda(\|f\|_{*}+\alpha_{0})$
$\leq f(-\lambda s)+\alpha_{0}\|\lambda s\|+\lambda(\alpha_{2}-\alpha_{0})$
$=f(-\lambda s)+\alpha_{0}\|\lambda s\|+(\alpha_{2}-\alpha_{0})\|\lambda s\|$
$=f(-\lambda s)+\alpha_{2}\|-\lambda s\|<0,$
where the second equality follows by the linearity of $f\in X^{*}$, the second inequality by the triangle inequality, the third inequality by the relation $f(-b)\leq|f(-b)|\leq\|f\|^{*}\|b\|$, and the fourth inequality by the fact that $\|b\|\leq~1$. Now, since $C_{\delta}\ =\mbox{cone}(\overline{S_{X}\cap C+\delta B_{X}})\subset\mbox{cone}(S_{X}\cap C+2\delta B_{X})$, it follows that for every $\alpha\in[\alpha_{1},\alpha_{0}]$ we have $f(x)+\alpha\|x\|<0<f(y)+\alpha\|y\|$ whenever $x\in-C_{\delta}\setminus\{0_{X}\}$ and $y\in\mbox{bd}(-K)\setminus\{0_{X}\}$. Furthermore, the inequality $f(x)+\alpha\|x\|<0$ for every $x\in-C_{\delta}\setminus\{0_{X}\}$ leads to $f\in(C_{\delta})^{a\#}_{+}$. Then, Theorem 3.2 (ii) $\Rightarrow$ (i) applies, concluding that the pair $(C_{\delta},K)$ satisfies the SSP. ∎
@@
[…]
@@
$0_{X}\not\in t\overline{\mbox{co}(C\cap S_{X})-\mbox{co}((\mbox{bd}(K_{t})\cap S_{X})\cup\{0_{X}\}))},\mbox{ for any }t>0.$
@@
[…]
@@ 4. Separation Theorems for Co-Radiant Sets
Then, $\text{(i)}\Rightarrow\text{(ii)}$. Moreover, if $\mathcal{C}$}} also admits a norm base, then $\text{(ii)}\Rightarrow\text{(i)}$.
@@
From now on, given $f\in X^{*}$ and $\alpha>0$, we consider the function $p_{f,\alpha}:X\rightarrow[0,+\infty)$ defined by $p_{f,\alpha}(x):=f(x)+\alpha\|x\|$, $\forall x\in X$. Since $f$ is linear and $\|\cdot\|$ is a norm, it follows that $p_{f,\alpha}$ is sublinear. The following lemma establishes some inequalities that describe the behavior of the level sets of $p_{f,\alpha}$.
@@
    For every $x\in X$ we have the inequalities $\|f\|_{*}\|x\|\leq p_{f,\alpha}(x)<2\|f\|_{*}\|x\|$.
@@
    $\rho f(x^{\prime})+\alpha\rho<-\rho(\alpha+\beta)+\alpha\rho=-\rho\beta.$
@@
Let $C\subset X$ be a cone, take $0<\delta<1$, $d>0$, and $y\in C_{\geq d}$. Then $y+\delta dB_{X}\subset(C_{\delta})_{\geq(1-\delta)d}$.
@@
[…]
@@
for every $y\in-C_{\delta}\setminus\{0_{X}\}$ and $y^{\prime}\in\mbox{bd}(-K)\setminus\{0_{X}\}$. Furthermore, $f\in{{\mathcal{C}}}^{\#}$ because $f\in(C_{\delta})^{\#}\subset C^{\#}$ and ${{\mathcal{C}}}\subset C\setminus\{0_{X}\}$. On the other hand, without loss of generality, we may assume that $f\in S_{X^{*}}$. Define $\lambda:=\inf\{f(y)-\alpha\|y\|:y\in{{\mathcal{C}}},\alpha\in(\alpha_{1},\alpha_{2})\}$, we will prove that $\lambda>0$ or, equivalently, that $\lambda=\sup\{p_{f,\alpha}(y):y\in-{{\mathcal{C}}},\alpha\in(\alpha_{1},\alpha_{2})\}<0$. The inequality $\lambda\leq 0$ is a direct consequence of (2) because ${{\mathcal{C}}}\subset-C_{\delta}\setminus\{0_{X}\}$. Now assume, contrary to our claim, that $\lambda=0$. For every $\alpha\in(\alpha_{1},\alpha_{2})$, we have $(f,\alpha)\in(C_{\delta})^{a\#}_{+}\subset C^{a\#}_{+}$, which implies $\alpha<1$ as $f$ is taken in $S_{X^{*}}$. Then $\alpha_{2}\leq 1$. However, without loss of generality, we may suppose that $\alpha_{2}<1$ because otherwise we can redefine $\alpha_{2}\in(\alpha_{1},1)$. Now, consider $\mu=\frac{1-\alpha_{2}}{2}\delta d_{{{\mathcal{C}}}}>0$. By the assumption $\lambda=0$, there exist $y_{0}\in-{{\mathcal{C}}}\subset-C_{\geq d_{{{\mathcal{C}}}}}$ and $\alpha_{0}\in(\alpha_{1},\alpha_{2})$ such that
$\mu<p_{f,\alpha_{0}}(y_{0})<0.$ (3)
@@
$\lambda<p_{f,\alpha}(y^{\prime}),\,\forall\alpha_{1}<\alpha<\alpha_{2},\,y^{\prime}\in\mbox{bd}(-{{\mathcal{K}}}(\eta)).$ (6)
To prove (6), we fix $\xi>0$ such that $\eta=\frac{\lambda}{\xi+I_{{{\mathcal{K}}}}}$ and define $\tau:=\lambda\frac{\frac{\xi}{2}+I_{{{\mathcal{K}}}}}{\xi+I_{{{\mathcal{K}}}}}<\lambda$. By Proposition 4.2, we have $I_{{{\mathcal{K}}}(\eta)}=\eta I_{{{\mathcal{K}}}}<(\frac{\xi}{2}+I_{{{\mathcal{K}}}})\eta=\tau$. Now pick an arbitrary ${\bar{y}}\in\mbox{bd}(-{{\mathcal{K}}}(\eta))$. If $\|{\bar{y}}\|\geq\tau$, then, by Lemma 3.7 (v), ${\bar{y}}\in\mbox{bd}(\mbox{cone}(-{{\mathcal{K}}}(\eta)))=\mbox{bd}(\mbox{cone}(-{{\mathcal{K}}}))=\mbox{bd}(-{{\mathcal{K}}})$. Now (2) applies and we have $\lambda<0<p_{f,\alpha}(\bar{y})$, for every $\alpha\in(\alpha_{1},\alpha_{2})$. To complete the proof, assume now that $\|{{\bar{y}}}\|<\tau$. Then Lemma 4.3 (iii) assures that $\lambda<-\tau\leq p_{f,\alpha}(\bar{y})$, for every $\alpha\in(\alpha_{1},\alpha_{2})$ finishing the proof of (6).
[…]
@@
[…]
@@
[…]
@@
    ${{\mathcal{K}}}_{3}:=\{x\in\mathbb{R}^{2}:f(x)-\alpha\|x\|_{2}\geq\lambda\}$, for any $f\in(\mathbb{R}^{2})^{*}$, $0<\alpha<\|f\|_{*}$, and $\lambda>0$, being $\|\cdot\|_{2}$ the Euclidean norm on $\mathbb{R}^{2}$.
Despite being uncommon for a co-radiant set to admit a norm base, what does occur is that any co-radiant set can be expressed as a union of a monotonically non-decreasing sequence (in the sense of inclusion) of co-radiant subsets with a norm base. Indeed, let ${{\mathcal{K}}}\subset X$ be a co-radiant set. Let us recall that, for every $r>0$, we define the set ${{\mathcal{K}}}_{r}:={{\mathcal{K}}}\cap rS_{X}$. Now, for every $n\in\mathbb{N}$ we define ${{\mathcal{K}}}^{n}:={{\mathcal{K}}}\cap\mbox{cone}({{\mathcal{K}}}_{n})$, if ${{\mathcal{K}}}_{n}\not$$\emptyset$, and ${{\mathcal{K}}}^{n}=\emptyset$, otherwise. The following lemma is easy to check.
@@
    If ${{\mathcal{K}}}_{n}\not$$\emptyset$, then ${{\mathcal{K}}}_{n}={{\mathcal{K}}}^{n}\cap nS_{X}$ and ${{\mathcal{K}}}_{n}$ is a norm-base for ${{\mathcal{K}}}^{n}$.
@@
Let $X$ be a normed space, and let $\mathcal{C}$}} and $\mathcal{K}$}} be co-radiant subsets of $X$ such that $d_{{{\mathcal{C}}}}d_{{{\mathcal{K}}}}>0$ and $\overline{co}({{\mathcal{C}}})\cap{{\mathcal{K}}}\not$$\varnothing$. Assume that there exists $n\in\mathbb{N}$ such that the pair $(\mbox{cone}({{\mathcal{C}}}),\mbox{cone}({{\mathcal{K}}}_{n}))$ satisfies the SSP. Then, there exist $\alpha_{1}$, $\alpha_{2}\in\mathbb{R}$ with $0<\alpha_{1}<\alpha_{2}$ and $f\in S_{X^{*}}$ such that:
@@
    $f(y^{\prime})-\alpha\|y^{\prime}\|<\lambda<f(y)-\alpha\|y\|,$